# Reinforced Diffuser for Red Teaming Large Vision-Language Models

### v1

@@ 5.3 Visualizations
Figure 3: Comparison of images generated by the pre-trained Stable Diffusion (left) and the Red Team Diffuser (right). Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
6 Limitations
The RTD framework demonstrates strong performance in eliciting harmful outputs from VLMs, but several limitations deserve attention. First, this study focuses exclusively on toxic continuation tasks. Broader challenges, such as misinformation propagation and ethical biases in complex multimodal scenarios, remain unexplored. Second, RTD’s high computational demands, including the use of 8 NVIDIA A100 GPUs, may hinder its adoption by researchers with limited resources. Finally, RTD poses ethical risks, as its methods could potentially be misused by adversaries.
7 Conclusion
In this work, we propose the Red Team Diffuser (RTD) framework to identify and exploit vulnerabilities in VLMs, specifically in toxic continuation tasks. Leveraging greedy prompt search and reinforcement fine-tuning, RTD generates images that effectively trigger toxic continuations, exposing the limitations of existing alignment mechanisms. Experimental results validate the effectiveness of RTD, showing significant increases in toxicity rates across both black-box and cross-model settings. For instance, RTD raises LLaVA’s toxicity rates by 10.69% on the original attack set and 8.91% on a hold-out set, while also demonstrating strong cross-model transferability, with increases of 5.1% on Gemini and 26.83% on Llama-3.2-11B-Vision-Instruct. Qualitative analysis further highlights RTD’s ability to subtly manipulate visual features, generating images that are more provocative and capable of inducing harmful continuations. These findings highlight critical gaps in current alignment strategies and emphasize the need for more robust and comprehensive mechanisms to enhance the safety of VLMs in real-world applications.

### v2

@@ 5.2. Cross-Model Transferability
Figure 3. Comparison of images generated by the general-purpose Stable Diffusion (left) and the Red Team Diffuser (right). Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
Table 5. CLIPScore comparison across different diffusers.
Stable Diffusion RTD (w/o) RTD (L) RTD (M) RTD (H)
CLIPScore 44.47 41.87 42.86 44.02 44.80
5.3. Visualizations
Figure 3 presents a qualitative comparison between images generated by the general-purpose Stable Diffusion model (left) and our Red Team Diffuser (RTD) (right). For RTD-generated images, the alignment reward weights $\lambda$ progressively increase from left to right, starting with a weight of 0 in the leftmost column and followed by 0.05, 0.1, and 0.2. Although all models receive the same image prompts, the resulting images diverge markedly in tone and visual implications. While the general-purpose Stable Diffusion consistently produces neutral and benign visuals. In contrast, RTD generates images that subtly evoke discomfort or toxicity-relevant cues. For instance, in the second row, RTD produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion. In the third row, where the prompt involves generating an image of a woman, RTD creates visuals that are noticeably more sexually suggestive and explicit. Similarly, in the fourth row, RTD introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations. The observed differences in visual elements and stylistic features underscore RTD’s ability to adapt visual semantics in ways that exploit the vulnerabilities of target VLMs. The comparison highlights how reinforced optimization—not merely prompt selection—enables subtle but impactful visual manipulation aligned with toxic continuation goals.
To complement these qualitative observations, we report CLIPScore (Hessel et al., 2021) as a quantitative measure of image-text semantic alignment. As shown in Table 5, all RTD variants achieve CLIPScores comparable to or higher than Stable Diffusion, indicating that the adversarial images remain semantically faithful to their prompts despite conveying more toxic or discomforting signals. Notably, RTD with the highest alignment reward ($\lambda=0.2$) attains a CLIPScore of 44.80, surpassing the baseline (44.47) and demonstrating that our stealth-aware reward mechanism enhances both visual regularity and the adversarial effectiveness of the generated images.
6. Limitations
While the Red Team Diffuser (RTD) framework shows strong performance and cross-model transferability, it has several limitations: 1. Task specificity. RTD is designed for toxic continuation tasks, and its effectiveness on other multimodal attacks, such as instruction-level jailbreaks, is yet to be explored. 2. Computational cost. The RL-based fine-tuning requires significant GPU resources, which may be impractical for resource-constrained users.
7. Conclusion
We present Red Team Diffuser (RTD), a novel framework for uncovering critical vulnerabilities in Vision-Language Models (VLMs) through semantically adversarial image generation. We formalize cross-modal toxicity amplification as a new attack surface, where visually coherent yet maliciously tuned images exacerbate harmful text continuations—an aspect largely overlooked by current alignment efforts. RTD combines greedy prompt search with reinforcement fine-tuning to generate images that reliably trigger toxic outputs. Experimental results validate its effectiveness: RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set and 8.91% on a hold-out set, while demonstrating strong cross-model transferability, achieving toxicity gains of 5.1% on Gemini and 26.83% on LLaMA-3.2-11B-Vision-Instruct. Qualitative analysis further illustrates how RTD manipulates visual cues to subtly induce toxic continuations. These findings expose substantial blind spots in existing VLM alignment strategies and underscore the need for more robust, multimodal-aware safety mechanisms. In future work, we aim to expand RTD to broader adversarial objectives and model families, offering deeper insights into VLM vulnerabilities and guiding the development of stronger alignment and defense strategies.